OpenStudy (anonymous):

Suppose you choose a marble from a bag containing 4 red marbles,2 whites marbles and 3 blue marbles. You return the first marble to the bag and then choose again. Find P (red and blue)

OpenStudy (mathmagician):

(4/9)*(3/9)=12/81=4/27